OpenStudy (anonymous):

Write the number 2011 using only the digit 4 and any of the operations of addition, subtraction, multiplication, division, exponentiation, taking a square root and factorial. You can use any number composed of the digit 4, even if it's decimal, so 44 and 44.44 are both allowed. You're also allowed to use brackets. Have fun!

OpenStudy (anonymous):

does 4(4^4) + 4(4^4) - (4!) - 4 - 4 - 4 - 4/4 count?

OpenStudy (yttrium):

\[4(444)+4^4-(4x4)-4-\frac{ 4 }{ 4 } = 2011\] I got this!

OpenStudy (anonymous):

@Yttrium You also got it but @sourwing is the first.. :))) Anyways, I'm a fan..

OpenStudy (anonymous):

:DD thanks

OpenStudy (yttrium):

another teaser?

OpenStudy (anonymous):

hold on a real quick :))

OpenStudy (anonymous):

How can you add eight 8's to get the number 1,000? (only using addition) @Yttrium That's an easy one :)

hartnn (hartnn):

888+88+8+8+8

OpenStudy (anonymous):

Wait, if you allow \(4/4\) then you could do \[ \underbrace{\frac 44+\frac 44+\dots +\frac 44}_{2011} \]That's lame...

OpenStudy (yttrium):

888+88+8+8+8 ugh

hartnn (hartnn):

yeah, there should be restrictions on number of 4s

OpenStudy (anonymous):

Haha. @Yttrium Can't you make it more quick? :)) Cheers for @hartnn

OpenStudy (yttrium):

net problems lol >.<

OpenStudy (the_fizicx99):

Wait but wouldn't 4/4 = 1 ? that would be the same thing as adding 1 + 1 + 1 ... till you get to 2011 >.> yeah that is lame tehe

OpenStudy (anonymous):

That's not the point, guys.. The challenge is to use different operations and symbols...

OpenStudy (yttrium):

tease us more

OpenStudy (the_fizicx99):

True, I like how you're only allowed to use 4's but you used 4/4 to get a 1 to subtract. Using exponent can be tricky to work with, especially since you were only able to use 4's. That was smart tbh >.<